OpenStudy (anonymous):

If the first equation is multiplied by 3 and then the equations are added, the result is _____. 3x + y = 3 x - 3y = -2 10x= 11 10x= 7 8x= 7

OpenStudy (anonymous):

lets multiply the 1st equation by 3 3(3x + y = 3) 9x + 3y = 9 9x + 3y = 9 x - 3y = -2 ----------add 10x = 7 any questions ?